Schrodinger's equation for a 3D cube
1
Expert's answer
2020-04-26T18:49:19-0400

The Schrödinger equation in 3D is


"-\frac{\hbar^2}{2m}\bigg(\frac{\partial^2\psi}{\partial x^2}+\frac{\partial^2\psi}{\partial y^2}+\frac{\partial^2\psi}{\partial z^2}\bigg)=E\psi(x,y,z)."

The wavefunction for a cubic box with side L is


"\psi(x,y,z)=A\text{ sin}\bigg(\frac{n_1\pi x}{L}\bigg)\text{ sin}\bigg(\frac{n_2\pi y}{L}\bigg)\text{ sin}\bigg(\frac{n_3\pi z}{L}\bigg)."

The normalization constant is


"A=\bigg(\frac{2}{L}\bigg)^{3/2}."